OpenStudy (anonymous):

Convert the mixed number, 1 3/4 , to a fraction. Reduce fraction if necessary

OpenStudy (asnaseer):

This site should help you find the answer: http://www.mathsisfun.com/mixed-fractions.html

OpenStudy (anonymous):

multiply 4 by 1 and add 3, because 1=4/4, and 4/4+3/4=7/4